A depth finder shows that the water in a certain place is 620 feet deep. The difference between d, the actual depth of the water, and the reading is │d-620│ and must be less than or equal to 0.05d. Find the minimum and maximum values of d, to the nearest tenth of a foot. {Please show work}
Answer by stanbon(75887)

│d-620│ and must be less than or equal to 0.05d. Find the minimum and maximum values of d, to the nearest tenth of a foot.
-------------------
Solve: |d-620] <= 0.05d
d-620 <= 0.05d and d-620 >=-0.05d
0.95d <= 620 and 1.05d >= 620
-----------------------
d <= 652.63 and d >= 590.48
-----------------------------
590.48 <= d <= 652.63
============================
Cheers,
Stan H.
============================